what is the prime factorization of the number 12600?
Answer by drk(1908)

There a few ways to attack this problem, here is one of them.
12600 / 2 = 6300
6300 / 2 = 3150
3150 / 2 = 1575
1575 / 3 = 525
525 / 3 = 175
175 / 5 = 35
35 / 5 = 7
7 / 7 = 1
The prime factorization is 2^3*3^2*5^2*7.